You have gathered the following information on your investments. What is the expected return on the portfolio?

Stock Number of Shares Price per Share Expected Return
F 210 $ 30 12.92 %
G 265 $ 16 9.55 %
H 205 $ 42 10.29 %

Multiple Choice

  • 11.47%

  • 10.92%

  • 12.44%

  • 11.96%

  • 10.99%
Answer #1
Stock Total value=(Number of shares*Price per share) Expected return
F (210*30)=6300 12.92%
G (265*16)=4240 9.55%
H (205*42)=8610 10.29%
Total value=$19150

Expected return of portfolio=Respective return*Respective weight

=(6300/19150*12.92)+(4240/19150*9.55)+(8610/19150*10.29)

=10.99%(Approx).